i have homework but i can't figure it out. please help i'll give you the problem to see if you can figure it out. i have the answer just not the actual work. please help. here it is.
2y + 7 = y
the answer is y = -7 but i can't get how to get to the answer Help!it seems so simple but i just can't get it




Answer by stanbon(75887)

2y + 7 = y
Subtract "y" from both sides of the equation, as follows:
2y + 7 -y = y-y
Then simplify both sides to get:
y+7 = 0
Now, subtract "7" from both sides to get:
y+7-7 = 0-7
y = -7
Cheers,
stan H.